# Configurations, Tessellations and Tone Networks

### v4

@@ XI Finale
SIMPLICIO – What is on the program for the concert tonight? I did not have a chance to check it out.
@@
SALVIATI – Good thing you got here eventually. We wouldn’t want to miss that! But Hagen has very little to sing in that last scene. I presume that the soprano will be singing Brünnhilde’s final monologue as she prepares to hurtle herself into the fire?
@@
SAGREDO – A diminished chord, not easy to get right. And these are the same notes that Gurnemanz, also a bass, sings, in the same key signature, four flats, on the black knight’s entry in Parsifal at the beginning of Act 3. Do you think there is some connection between Gurnemanz and Hagen?
SALVIATI – You are always spotting these connections! Well, it’s more likely that Wagner was just recycling some of his ideas. Maybe he was running short of material. But, Simplicio, don’t you have to prepare yourself a bit and get ready in the rehearsal rooms downstairs?
@@
SALVIATI – Well, that is an interesting story. What they told me was the following. In an earlier draft of their paper, they introduced their new approach to the tonnetz as a configuration in a purely triadic context and then analyzed some passages in Tristan and Parsifal using a reductive approach – that is, using a triadic tonnetz-based analysis to study various tetradic harmonies appearing in the operas. And they made some interesting comparisons with Chopin and Tchaikovsky, who used some of the same chords .  .  .
SIMPLICIO – Please, go on.
SALVIATI – Of course. But you see, Simplicio, a couple of months ago, when they were presenting their work at the University of Southern North Dakota, someone in the audience complained that they needed to justify this reductive approach. Also, in the same talk, they made some remarks that suggested (I am not sure they really meant this) that in their opinion all previous attempts over the past few decades to find a mathematically satisfactory tonnetz theory of tetradic harmonies had been unsuccessful. The audience seemed to be unhappy with this, or so it appeared, so Bolando and Ugopoli felt that the only way around this impasse – this is what they told me – would be to create a new theory of tetradic harmony based on a tonnetz that was associated with a configuration.
@@
SALVIATI – Well, you might have thought so, but the situation was not as simple as that. For sure, a lot of good work has been undertaken by music theorists on the relations between dominant seventh chords and half-diminished sevenths – Bolando and Ugopoli for some reason insist on calling them minor sixths – but the music theorists would typically take a ‘maximalist’ approach, looking at all possible transitions (after all, Wagner uses a lot of them), throwing in various other chords as well, for good measure. Just look at Cohn’s book, in the section called, ‘Circumnavigating the Tristan-Genus Universe’. The writing is brilliant, I have to admit, but as a basis for his analysis he uses the Douthett 4-Cube Trio, which is an enormous graph involving twelve dominant seventh chords, twelve minor sixth chords, twelve minor seventh chords, six French sixths,   .  .  .
@@
SAGREDO – You are saying that such a graph, although demonstrably useful, at least in some ways, for descriptive purposes, is unlikely to be recognizable as a mathematical object?
@@
SIMPLICIO – But what exactly do you mean by ‘recognizable as a mathematical object’?
@@
SAGREDO – Are you saying that there are aesthetic elements involved here? Is it a question of mathematical good taste, whatever that means?
@@
SALVIATI – Well, Simplicio, who knows exactly how they got there. Often, hitting on the right idea is just a matter of good luck. But from what I understand, they had been looking at the relations between triads when exactly two notes could change in the transition from a major chord to a minor chord. This led to a certain bipartite graph with girth four that could be unfolded into an Archimedean tessellation of the plane, the famous one of Kepler.
SIMPLICIO – Girth four means that there are tetracycles and that there is no configuration.
SALVIATI – That is true, but the tessellation can be regarded as a species of tonnetz and it is also a geometrical object in its own right. And it gives rise to interesting cycles.
SAGREDO – So maybe we should not be so dogmatic about what is a tonnetz and what is not. Perhaps we should simply be patient and accept what the mathematics tells us.
@@
SIMPLICIO (nodding in agreement) – Yes, of course, but you were saying  .  .  .
@@
SALVIATI – Yes, they added in the constraint that the ‘dissonant’ note in a given tetrachord should be preserved under the transition. This can be done consistently. The flatted seventh in the dominant seventh chord is preserved in the resulting minor sixth, and the sixth of a minor sixth chord is preserved in the resulting dominant seventh. The result, miraculously, is a Levi graph, with no tetracycles, that represents another configuration $\{12_{3}\}$. This time, the D228. So that is how they got their mathematical object.
@@
SALVIATI – Well, yes, of course, and that is why we came to tonight’s performance, so we could hear with our own ears the famous octacycle of Bolando and Ugopoli being played as Brünnhilde approaches her immolation.
@@ Acknowledgements.
LPH wishes to acknowledge E. Chew, M. Gotham, K. Rietsch and other members of the Music and Acoustics Research Centre (MARC) at King’s College London for stimulating discussions. The authors are grateful for comments made by participants at a MARC seminar in May 2025, and at a colloquium of the Center for Theoretical Physics of the Polish Academy of Sciences in December 2025, where a drafts of this work have been presented. We are grateful for the comments and suggestions of two anonymous reviewers, which have led to many improvements. The authors also wish to thank Abdulla Alazemi, J. Armstrong, D. M. Blasius, L. Masiero, R. Morales, P. Nurowski, S. Salamon and others for helpful discussions and correspondence.

### v5

@@ XI Finale
SIMPLICIO – What is on the program for the concert tonight?
@@
SALVIATI – Good thing you got here eventually. We wouldn’t want to miss that! But Hagen has very little to sing in that last scene. I presume that the soprano will be singing Brünnhilde’s great monologue as she prepares to hurtle herself into the fire?
@@
SAGREDO – A diminished chord, not easy to get right. And these are the same notes that Gurnemanz, also a bass, sings, in the same key signature, four flats, in Parsifal at the beginning of Act 3, when Parsifal returns as the black knight. Do you think there is some connection between Gurnemanz and Hagen?
SALVIATI – You are always spotting these connections! Well, it’s more likely that Wagner was just recycling some of his ideas. Maybe he was running short of material. But, Simplicio, don’t you have to prepare yourself and get ready in the rehearsal rooms downstairs?
@@
SALVIATI – Well, that is an interesting story. In an earlier draft of their paper, they introduced their new approach to the tonnetz as a configuration in a purely triadic context. They analyzed passages in Tristan and Parsifal using a reductive approach – that is, using a triadic tonnetz to study tetradic harmonies in the operas. And they made comparisons with works of Chopin and Tchaikovsky, who used some of the same chords. However, not long ago, when they were presenting their work at the University of Southern North Dakota, someone in the audience argued that they needed to justify this reductive approach.
SAGREDO – Not unreasonable, I suppose.
SALVIATI – Also, in the same talk, Bolando and Ugopoli made some remarks to the effect that in their opinion all previous attempts to find a mathematically satisfactory tonnetz theory of tetradic harmonies had been unsuccessful. The audience seemed to be unhappy with this claim, or so it appeared, so Bolando and Ugopoli felt that the only way to resolve the matter – this is what they told me – would be to create a new theory of tetradic harmony based on a tonnetz that was associated with a configuration.
@@
SALVIATI – Well, you might have thought so, but the situation was not as simple as that. For sure, a lot of good work has been undertaken by music theorists on the relations between dominant seventh chords and half-diminished sevenths – or minor sixths – but the music theorists would typically take a ‘maximalist’ approach, looking at all possible transitions between these chords (after all, Wagner uses a lot of them), throwing in various other chords as well, for good measure. Just look at Cohn’s book, in the section called, ‘Circumnavigating the Tristan-Genus Universe’. The writing is brilliant, I have to admit, but as a basis for his analysis he uses the Douthett 4-Cube Trio, which is an enormous graph involving twelve dominant seventh chords, twelve minor sixth chords, twelve minor seventh chords, six French sixths,   .  .  .
@@
SAGREDO – You are saying that such a graph, though demonstrably useful, at least in some ways, for descriptive purposes, is unlikely to be recognizable as a mathematical object?
@@
SIMPLICIO – But what exactly do you mean, in the present context, by ‘recognizable as a mathematical object’?
@@
SAGREDO – Are you saying that there are aesthetic elements involved here? Is it a question of ‘mathematical good taste’, whatever that means?
@@
SALVIATI – Well, Simplicio, who knows exactly how they got there. Often, hitting on the right idea is just a matter of good luck or intuition. But from what I understand, they had been looking at the relations between triads when exactly two notes could change in the transition from a major chord to a minor chord. This led to a certain bipartite graph of girth four and degree three that could be unfolded into an Archimedean tessellation of the plane, the famous one of Kepler involving dodecagons, hexagons and squares.
SIMPLICIO – Girth four means that there are tetracycles – these are the squares – and that there is no configuration.
SALVIATI – That is true, but the tessellation can be regarded as a species of tonnetz and it is also a geometric object in its own right. And it gives rise to interesting cycles.
SAGREDO – So, maybe we should not be so dogmatic about what is a tonnetz and what is not. Perhaps we should simply be patient and accept what the mathematics tells us.
@@
SIMPLICIO (nodding in agreement) – Yes, of course, preconceptions lead to misconceptions. But you were saying  .  .  .
@@
SALVIATI – Yes, they added in the constraint that the ‘dissonant’ note in a given tetrachord should be preserved under the transition. This can be done consistently. The flatted seventh in the dominant seventh chord is preserved in the resulting minor sixth, and the sixth of a minor sixth chord is preserved in the resulting dominant seventh. The result, miraculously, is a Levi graph, with no tetracycles, that represents another configuration $\{12_{3}\}$. This time, the D228. And that is how they got their mathematical object.
@@
SALVIATI – Well, yes, of course, and that is why we came to tonight’s performance, so we could hear with our own ears the octacycle of Bolando and Ugopoli being played as Brünnhilde approaches her immolation.
@@ Acknowledgements.
LPH wishes to acknowledge E. Chew, M. Gotham, K. Rietsch and other members of the Music and Acoustics Research Centre (MARC) at King’s College London for stimulating discussions. The authors are grateful for comments made by participants at a MARC seminar in May 2025, and at a colloquium of the Center for Theoretical Physics of the Polish Academy of Sciences in December 2025, where drafts of this work have been presented. We are grateful for the comments and suggestions of two anonymous reviewers, which have led to many improvements. The authors also wish to thank Abdulla Alazemi, J. Armstrong, D. M. Blasius, J. Forth, L. Masiero, R. Morales, P. Nurowski, S. Salamon and others for helpful discussions and correspondence.
APPENDIX: Cycle Count Tables
In this appendix we record the numbers of cycles of various types for the Eulerian tonnetz, the Tristan tonnetz, the Pentatonic tonnetz, the Twelve-tone tonnetz, and the Archimedean tonnetz. The rows are labelled by cycle length and the columns are labelled by the numbers of crossings (the $p$-numbers). Each entry in the table shows the number of cycles of a given length with the given number of crossings. Totals are tabulated for each row and each column. For example, in the case of the Eulerian tonnetz we observe that there are twelve 2pp-hexacycles, four 3pp-hexacycles, and a total of sixteen hexacycles altogether, and that the Eulerian tonnetz admits a grand total of 5409 cycles.
Cycle Count for Eulerian Tonnetz
0 11 22 33 44 55 66 77 88 99 $10$ $11$ Total
66 0 0 $12$ 44 0 0 0 0 0 0 0 0 $16$
88 0 $12$ 0 $24$ 33 0 0 0 0 0 0 0 $39$
$10$ 0 0 $12$ $48$ $60$ 0 0 0 0 0 0 0 $120$
$12$ 0 0 $30$ $72$ $78$ $48$ 0 0 0 0 0 0 $228$
$14$ 0 0 $24$ $24$ $228$ $192$ $120$ $48$ 0 0 0 0 $636$
$16$ 0 0 0 $72$ $168$ $324$ $186$ $156$ 0 0 0 0 $906$
$18$ 0 $12$ $12$ $36$ $108$ $324$ $468$ $348$ $192$ 0 0 0 $1500$
2020 0 0 0 $48$ $42$ $252$ $192$ $312$ $174$ $132$ $30$ 0 $1182$
$22$ 0 0 $12$ 0 $48$ $48$ $84$ $108$ $228$ $144$ $36$ $12$ $720$
$24$ 11 0 0 0 0 0 66 $36$ 33 44 0 $12$ $62$
Total 11 $24$ $102$ $328$ $735$ $1188$ $1056$ $1008$ $597$ $280$ $66$ $24$ $5409$
Table 5: Cycle count for the Eulerian tonnetz. The rows are labelled by cycle length and the columns are labelled by the number of $p$-crossings.
Cycle Count for Tristan tonnetz
0 11 22 33 44 55 66 77 88 99 $10$ $11$ Total
66 0 $12$ $12$ 0 0 0 0 0 0 0 0 0 $24$
88 0 0 0 $12$ 33 0 0 0 0 0 0 0 $15$
$10$ 0 0 $12$ $24$ $60$ $24$ 0 0 0 0 0 0 $120$
$12$ 0 0 0 $40$ $78$ $96$ 22 0 0 0 0 0 $216$
$14$ 0 0 0 $72$ $120$ $240$ $192$ 0 0 0 0 0 $624$
$16$ 0 0 $42$ $12$ $144$ $228$ $276$ $144$ 0 0 0 0 $846$
$18$ 0 0 $12$ $72$ $108$ $192$ $364$ $444$ $192$ $40$ 0 0 $1424$
2020 0 $12$ 0 $24$ $54$ $84$ $186$ $372$ $282$ $168$ 0 0 $1182$
$22$ 0 0 $12$ $12$ $12$ $48$ $96$ $120$ $168$ $120$ $24$ $12$ $624$
$24$ 11 0 0 0 0 0 0 0 33 $16$ $18$ $12$ $50$
Total 11 $24$ $90$ $268$ $579$ $912$ $1116$ $1080$ $645$ $344$ $42$ $24$ $5125$
Table 6: Cycle count for the Tristan tonnetz. The rows are labelled by cycle length and the columns are labelled by the number of $p$-crossings.
Cycle Count for Pentatonic Tonnetz
0 11 22 33 44 55 66 77 88 99 $10$ Total
66 0 55 $10$ 55 0 0 0 0 0 0 0 2020
88 0 0 $10$ 2020 0 0 0 0 0 0 0 $30$
$10$ 0 55 2020 $45$ $50$ $12$ 0 0 0 0 0 $132$
$12$ 0 55 55 $30$ $60$ $45$ 55 0 0 0 0 $150$
$14$ 0 0 2020 $40$ $120$ $125$ $110$ 55 0 0 0 $420$
$16$ 0 55 0 $25$ $60$ $105$ $80$ $85$ 0 0 0 $360$
$18$ 0 0 55 2020 $30$ $40$ $95$ $90$ $30$ $10$ 0 $320$
2020 11 0 0 0 0 77 55 0 55 55 11 $24$
Total 11 2020 $70$ $185$ $320$ $334$ $295$ $180$ $35$ $15$ 11 $1456$
Table 7: Cycle count for the pentatonic tonnetz. The rows are labelled by cycle length and the columns are labelled by the number of $p$-crossings.
Cycle Count for Twelve-Tone Tonnetz
0 11 22 33 44 55 66 77 88 99 $10$ $11$ $12$ $13$ $14$ Total
88 0 55 $30$ $45$ $10$ 0 0 0 0 0 0 0 0 0 0 $90$
$10$ 0 55 55 $30$ $25$ 77 0 0 0 0 0 0 0 0 0 $72$
$12$ 0 0 $10$ $70$ $135$ $80$ 55 0 0 0 0 0 0 0 0 $300$
$14$ 0 55 $40$ $105$ $295$ $365$ $255$ $15$ 0 0 0 0 0 0 0 $1080$
$16$ 0 0 $15$ $85$ $280$ $465$ $530$ $230$ $15$ 0 0 0 0 0 0 $1620$
$18$ 0 55 55 $80$ $350$ $845$ $1085$ $1130$ $320$ 2020 0 0 0 0 0 $3840$
2020 0 0 $35$ $135$ $325$ $1007$ $1685$ $2080$ $1655$ $580$ $22$ 0 0 0 0 $7524$
$22$ 0 55 0 $55$ $245$ $575$ $1415$ $2035$ $2070$ $1410$ $460$ $10$ 0 0 0 $8280$
$24$ 0 55 0 $45$ $155$ $535$ $855$ $1685$ $2345$ $2135$ $1335$ $345$ $10$ 0 0 $9450$
$26$ 0 0 $25$ $25$ $85$ $245$ $615$ $895$ $1390$ $1610$ $1410$ $965$ $285$ $10$ 0 $7560$
$28$ 0 0 0 $25$ 55 2020 $105$ $140$ $150$ $255$ $305$ $200$ $170$ $60$ 55 $1440$
$30$ 11 0 0 0 0 77 0 $15$ 2020 2020 $16$ $10$ $30$ 2020 55 $144$
Total 11 $30$ $165$ $700$ $1910$ $4151$ $6550$ $8225$ $7965$ $6030$ $3548$ $1530$ $495$ $90$ $10$ $41400$
Table 8: Cycle count for twelve-tone Tonnetz. The rows are labelled by cycle length and the columns are labelled by the number of $p$-crossings.
Cycle Count for the Archimedean Tonnetz
0 11 22 33 44 55 Total
44 0 0 33 0 0 0 33
66 0 66 $12$ 22 0 0 2020
88 0 66 0 $18$ 0 0 $24$
$10$ 0 0 $12$ $18$ $12$ 66 $48$
$12$ 11 0 0 22 33 66 $12$
Total 11 $12$ $27$ $40$ $15$ $12$ $107$
Table 9: Cycle count for the Archimedean tonnetz. The rows are labelled by cycle length and the columns are labelled by the number of $p$-crossings.